OpenStudy (anonymous):

Another question I need help with Write an equation in slope intercept form of the line that passes through (4, 1) and (5, -1). y = -4x + 1 y = -1/2 x + 9 y = 4x + 1 y = -2x + 9

OpenStudy (kira_yamato):

dy/dx = (-1 - 1) / (5 - 4) = -2 y = -2x + C When x = 4, y = 1, 1 = -2(4) + C C = 9 y = -2x + 9